Can you please walk me through the solution.... I have seen examples, but I missed class for 2 days & I am very lost at this point.
1.) y/4=5/6
2.) 6m/5=-6

Answer by ankor@dixie-net.com(12696)

1.) y/4 = 5/6
:
It becomes easy if you eliminate the denominators, mult eq by a common denomiator, 12 would be a good one, you have:
12(y/4) = 12(5/6)
:
Dividing the 4 and the 6 into 12 you have:
3y = 2(5)
:
3y = 10
:
Divide both sides by 3 and you have:
y = 10/3
:
:
2.) 6m/5=-6
:
mult equation by 5 to get rid of the denominator:
5(6m/5) = 5(-6)
:
6m = -30
:
m = -30/6
:
m = -5